OpenStudy (anonymous):

what are the zeroes of the function f(x)= 3x^2+9x+6 / 3x+3

OpenStudy (anonymous):

u just need to find zeroes of num

OpenStudy (anonymous):

\[(3x^2+9x+6)=(x+2)(3x+3)\] equating this to zero will give you zeroes

OpenStudy (anonymous):

would the zeroes be -2, -3?

OpenStudy (anonymous):

(x+2)(3x+3)=0 x=-2, -1, not sure tho :)

OpenStudy (anonymous):

im sorry i mean -2

OpenStudy (anonymous):

Okay thank you!!

OpenStudy (anonymous):

yeah x=-2 and -1

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

you need to leave the variable x by it'self

OpenStudy (anonymous):

okay thanks!

hero (hero):

Erica when you factored the trinomial, you forgot to rewrite the fraction as \[\frac{((x+2)(3x+3)}{(3x+3)}\] Had you done that, you would have realized that \(\large\frac{(3x+3)}{(3x+3)}\) cancels leaving you with x + 2 only, which means we would only need the zero for (x+2) which is x = -2

hero (hero):

@erica.d

hero (hero):

Especially since x = -1 is undefined

OpenStudy (anonymous):

@Hero thanks !! i just thought about the factors because of what value of numerator will be zero, hence i will have zero of that function ...